OpenStudy (anonymous):

Evaluate 5|7 - 9| - 2.

OpenStudy (anonymous):

Is that 5/7?

OpenStudy (anonymous):

5|7-9| - 2 = 5|-2| - 2 = 5(2) - 2 = 10-2 = 8. Reference the efinition of modulus: \[|x| = x\] if x>0, and \[|x| = -x\] if x<0 and obviously \[|0| = 0\] this is the definition of the norm on \[\mathbb{R}.\]